OpenStudy (anonymous):

x^5y^3z^3 ------ xy^-3z^-1

OpenStudy (goformit100):

@jackellyn

OpenStudy (anonymous):

wut?

OpenStudy (jackellyn):

x exponent 5-1=4 y exponent 3-(-3) = 6 z exp 3 -(-1) = 4 x^4y^6z^4

OpenStudy (jackellyn):

@imwhite remember than when you are dividing exponents you are subtracting them.

OpenStudy (anonymous):

ooohh, ok. thank you:)